### v1

@@ 5.1 Continuity condition on utilities
We first show that improved regret bounds are possible with a continuity condition on the player utilities. For intuition, the example in Table 1 gave a “hard” example resulting in linear regret in Theorem 3.1 and the related example in Table 3 resulted in $\Omega(T^{2/3})$ regret in Theorem 4.6. These examples relied on two outcomes with nearly identical utilities for the follower having significantly different utilities for leader, which could be viewed as a violation of continuity. This suggests that if arms that are sufficiently different for the leader were also sufficiently different for the follower, then it might be possible to beat the regret lower bound from Theorem 3.1 and Theorem 4.6.
@@
For example, when the two players have the same utilities (i.e., $v^{1}=v^{2}$), then $L^{*}=1$. More generally, our continuity condition captures the extent to which players agree on which outcomes are different from each other (a more detailed discussion is given in Section E.1). Returning to the examples in Tables 1, 3, the “hard” instances yielding linear regret for the original Stackelberg benchmarks (Theorem 3.1; Table 1) have $L^{*}=\Theta(T^{-1/2})$ and the corresponding “hard” instances for $T^{2/3}$ regret for the $\gamma$-tolerant benchmarks (Theorem 4.6; Table 3) require that $L^{*}=\Theta(T^{-1/3})$; in contrast, we focus on utility functions where $L^{*}$ is a constant.
When $L^{*}$ is bounded, we show that it is possible for both players to achieve $\tilde{O}(\sqrt{T})$ regret even with respect to the original Stackelberg benchmarks in the strongly decentralized setting. The follower can run any algorithm $\texttt{ALG}_{2}$ with sufficiently low high-probability anytime regret (e.g., ActiveArmElimination as in Proposition 7.1 or UCB as in Proposition 7.3). We construct another UCB-based algorithm LipschitzUCB (Algorithm 4) for the leader, which expands the confidence sets based on the Lipschitz constant $L^{*}$.
@@
The algorithm $\texttt{ALG}_{1}=\texttt{LipschitzUCB}(L,C)$ operates in the strongly decentralized setting, taking as inputs parameters $L$ and $C$. (The parameter $L$ is intended to be an upper bound on the Lipschitz constant $L^{*}$, and the parameter $C^{\prime}$ is intended to be such that $\texttt{ALG}_{2}$ satisfies anytime regret bound $h(t,T,\mathcal{B})=\sqrt{Ct\log T}$, where $C=C^{\prime}\cdot\sqrt{|\mathcal{B}|}$ for a constant $C^{\prime}$.) For each arm $a\in\mathcal{A}$, the algorithm computes UCB estimates $v^{\text{UCB}}_{1}(a)$ of the quantity $\max_{b\in\mathcal{B}}v_{1}(a,b)$ using the high-probability anytime regret bounds of $\texttt{ALG}_{2}$ as well as the upper bound on the Lipschitz constant. The algorithm then chooses the arm $a_{t}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a)$.
Initialize $\hat{v}_{1}(a)=1$ for $a\in\mathcal{A}$.   // Initialize empirical means for $\max_{b\in\mathcal{B}}v_{1}(a,b)$.
Initialize $v^{\text{UCB}}_{1}(a)=1$ for $a\in\mathcal{A}$.   // Initialize UCB for $\max_{b\in\mathcal{B}}v_{1}(a,b)$
@@
       Set $S(a):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},r)\in H\text{ s.t. }a=a_{t^{\prime}}\right\}$   // Observed rewards
@@
             $\hat{v}_{1}(a)\leftarrow(\sum_{r\in S(a)}r)/|S(a)|$   // Empirical mean
             $\alpha(a)\leftarrow\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{|S(a)|}}+C\cdot L\cdot\frac{\sqrt{\log T}}{\sqrt{|S(a)|}}$   // confidence bound width
@@ 3             $v^{\text{UCB}}_{1}(a)\leftarrow\min(1,\hat{v}_{1}(a)+\alpha(a))$
Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\left(v^{\text{UCB}}_{1}(a)\right)$.   // arm with max upper confidence bound
@@
[…]
@@
The intuition is the continuity conditions imply that small errors by the follower translate into bounded suboptimality for the leader (and vice versa); moreover, the high-probability anytime regret requirements bound the follower’s errors. Together, these properties guarantee that the leader’s empirical mean $\hat{v}_{1}(a)$ for each arm $a\in\mathcal{A}$ is close to the mean reward $v_{1}(a,b^{*}(a))$ that they would receive if the follower best-responded: in more detail, the main lemma (Lemma E.2) is that the empirical mean $\hat{v}_{1}(a)$ is at least $v_{1}(a,b^{*}(a))-\Theta(L\sqrt{\log T}/\sqrt{n_{t}(a)})$, where $n_{t}(a)$ is the number of times that arm $a$ has been pulled prior to time step $t$. Using Lemma E.2 to bound the suboptimality of the leader’s choice of actions $a_{t}\in\mathcal{A}$ and using the anytime regret requirements to bound the follower’s suboptimality, we can bound both the leader’s regret and the follower’s regret. We defer the full proof to Section E.2. ∎
@@ 5.2 Weaker benchmark
For the self-tolerant benchmarks, we show it is possible to achieve $\tilde{O}(\sqrt{T|\mathcal{A}||\mathcal{B}|})$ regret for both players (Theorem 5.3), which outperforms the $T^{2/3}$ lower bound for the stronger benchmark shown in Theorem 4.6. To demonstrate this is feasible, we construct a specific pair of algorithms (in the weakly decentralized setting where the leader can observe the follower’s actions), that achieve a $O(\sqrt{T})$ regret upper bound. For the follower, we take the algorithm $\texttt{ALG}_{2}$ to be ActiveArmElimination (Algorithm 7), which cycles through phases of exploration, after which all sufficiently suboptimal arms are eliminated. For the leader, we construct a UCB-based algorithm PhasedUCB (Algorithm 5) which constructs confidence bounds for every pair of actions $(a,b)$.
@@
The algorithm $\texttt{ALG}_{1}=\texttt{PhasedUCB}(M_{1},\ldots,M_{P})$ operates in the weakly decentralized setting, taking as input the parameters $M_{1},\ldots,M_{P}\geq 0$. (The parameter $M_{i}$ is intended to capture the number of times that an arm is pulled in phase $i$ by the instantiation of ActiveArmElimination specified by $\texttt{ALG}_{2}$.) The algorithm $\texttt{ALG}_{1}$ computes UCB estimates $v^{\text{UCB}}_{1}(a,b)$ for $v_{1}(a,b)$, computes the set of active arms $\mathcal{B}^{\prime}(a)$ in the previous phase of $\texttt{ALG}_{2}$’s instantiation of ActiveArmElimination for each arm $a$ (ComputeActiveArms, Algorithm 6), and chooses the arm with maximum UCB: $a_{t}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a,b)$. ComputeActiveArms computes the active arms $\mathcal{B}^{\prime}(a)$ by iterating through $H$ and keeping track of whenever a new phase is entered using the parameters $M_{1},\ldots,M_{P}$.
Let $\hat{v}_{1}(a,b)=0$ for $a\in\mathcal{A}$ and $b\in\mathcal{B}$// initialize empirical mean of $v_{1}(a,b)$
Let $v^{\text{UCB}}_{1}(a,b)=1$ for $a\in\mathcal{A}$ and $b\in\mathcal{B}$.   // initialize UCB for $v_{1}(a,b)$
Let $\mathcal{B}^{\prime}(a)=\texttt{ComputeActiveArms}(M_{1},\ldots,M_{P},H)$.   // active arms in previous phase for $\texttt{ALG}_{2}$
@@
             Set $S(a,b):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},b_{t^{\prime}},r)\in H\text{ s.t. }a=a_{t^{\prime}},b=b_{t^{\prime}}\right\}$   // observed rewards
@@
                   $\hat{v}_{1}(a,b)\leftarrow(\sum_{r\in S(a,b)}r)/|S(a,b)|$   // compute empirical mean
                   $\alpha(a,b):=10\cdot\sqrt{\frac{\log T}{|S(a,b)|}}$   // confidence bound width
                   $v^{\text{UCB}}_{1}(a,b)\leftarrow\min\left(1,\hat{v}_{1}(a,b)+\alpha(a,b)\right)$   // compute UCB
@@
Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}\left(v^{\text{UCB}}_{1}(a,b)\right)$.   // arm with max upper confidence bound for any valid $b$
@@
Initialize $s^{\prime}(a)=0$ for $a\in\mathcal{A}$.   // Index of the last completed phase for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $t^{\prime}(a)=1$ for $a\in\mathcal{A}$.   // Time step marking beginning of phase $s^{\prime}+1$ for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $\mathcal{B}^{\prime}(a)=\mathcal{B}$.   // Active arms in phase $s^{\prime}$ for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $newphase_{a}=\text{False}$ for $a\in\mathcal{A}$.   // Boolean for first time step in phase for $\texttt{ALG}_{2}$ on $a$.
@@
Suppose that for each $a\in\mathcal{A}$, the algorithm $\texttt{ALG}_{2}$ runs a separate instantiation of ActiveArmElimination with parameters $M_{1},\ldots,M_{P}$ (where $M_{i}=\Theta(\log T\cdot 2^{2i})$ denotes the number of times that each arm is pulled in phase $i$). Let $\texttt{ALG}_{1}=\texttt{PhasedUCB}(M_{1},\ldots,M_{P})$. Then it holds that the regret with respect to the self-$\gamma$-tolerant benchmarks $\beta_{1}^{\text{self-tol}}$ and $\beta_{2}^{\text{self-tol}}$ is bounded as:
@@
The intuition is that the benchmark allows the leader to choose any $a\in\mathcal{A}_{\epsilon}$. The definition of $\mathcal{A}_{\epsilon}$ means that the leader can take an optimistic perspective on the follower’s choice of action $\mathcal{B}_{\epsilon}(a)$ (and not have to prepare for the worst-case action in $\mathcal{B}_{\epsilon}(a)$). This optimistic perspective surfaces in PhasedUCB in terms of how the leader evaluates an action $a$ based on the maximum UCB $\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a,b)$ across all active arms $b\in\mathcal{B}^{\prime}(a)$. To analyze this pair of algorithms, we show a bound $\epsilon_{t}$ for each time step $t$ such that $a_{t}$ is an $\epsilon_{t}$-best-response for the leader and $b_{t}$ is an $\epsilon_{t}$-best-response for the follower: the main lemma (Lemma E.6) shows that we can set $\epsilon_{t}$ to be $\Theta(\sqrt{|\mathcal{B}\cdot\log T/n_{t}(a_{t})})$ where $n_{t}(a_{t})$ is the number of times that arm $a_{t}$ has been pulled prior to time step $t$. The full proof is deferred to Section E.3. ∎
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be any action sets such that $|\mathcal{A}|\geq 2$ and $|\mathcal{B}|\geq 2$. Consider any algorithms $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$ which operate in either the strongly decentralized setting or weakly decentralized setting. There exists an instance *$\mathcal{I}^{*}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ such that at least one of the players incurs $\Omega(\sqrt{T\cdot(\left|\mathcal{A}\right|-1)\cdot\left|\mathcal{B}\right|})$ regret with respect to the self-$\gamma$-tolerant benchmarks $\beta_{1}^{\text{self-tol}}$ and $\beta_{2}^{\text{self-tol}}$, that is: $\max(R_{1}(T;\mathcal{I}^{*}),R_{2}(T;\mathcal{I}^{*}))=\Omega(\sqrt{T\cdot(\left|\mathcal{A}\right|-1)\cdot\left|\mathcal{B}\right|})$.
@@
6 Discussion of Benchmark Parameters: The Maximum Tolerance and the Regularizer

### v2

@@ 5.1 Continuity condition on utilities
We first show that improved regret bounds are possible with a continuity condition on the player utilities. For intuition, the example in Table 1 gave a “hard” example resulting in linear regret in Theorem 3.1 and the related example in Table 3 resulted in $\Omega(T^{2/3})$ regret in Theorem 4.6. These examples relied on two outcomes with nearly identical utilities for the follower having significantly different utilities for the leader, which could be viewed as a violation of continuity. This suggests that if arms that are sufficiently different for the leader were also sufficiently different for the follower, then it might be possible to beat the regret lower bound from Theorem 3.1 and Theorem 4.6.
@@
For example, when the two players have the same utilities (i.e., $v_{1}=v_{2}$), then $L^{*}=1$. More generally, our continuity condition captures the extent to which players agree on which outcomes are different from each other (a more detailed discussion is given in Section D.1). Returning to the examples in Tables 1, 3, the “hard” instances yielding linear regret for the original Stackelberg benchmarks (Theorem 3.1; Table 1) have $L^{*}=\Theta(T^{-1/2})$ and the corresponding “hard” instances for $T^{2/3}$ regret for the $\gamma$-tolerant benchmarks (Theorem 4.6; Table 3) require that $L^{*}=\Theta(T^{-1/3})$; in contrast, we focus on utility functions where $L^{*}$ is a constant.
When $L^{*}$ is bounded, we show that it is possible for both players to achieve $\tilde{O}(\sqrt{T})$ regret even with respect to the original Stackelberg benchmarks. The follower can run any algorithm $\texttt{ALG}_{2}$ with sufficiently low high-probability anytime regret (e.g., ActiveArmElimination as in Proposition 7.1 or UCB as in Proposition 7.3). We construct another UCB-based algorithm LipschitzUCB (Algorithm 4) for the leader, which expands the confidence sets based on the Lipschitz constant $L^{*}$.
@@
The algorithm $\texttt{ALG}_{1}=\texttt{LipschitzUCB}(L,C)$ takes as inputs parameters $L$ and $C$. (The parameter $L$ is intended to be an upper bound on the Lipschitz constant $L^{*}$, and the parameter $C^{\prime}$ is intended to be such that $\texttt{ALG}_{2}$ satisfies anytime regret bound $h(t,T,\mathcal{B})=\sqrt{Ct\log T}$, where $C=C^{\prime}\cdot\sqrt{|\mathcal{B}|}$ for a constant $C^{\prime}$.) For each arm $a\in\mathcal{A}$, the algorithm computes UCB estimates $v^{\text{UCB}}_{1}(a)$ of the quantity $\max_{b\in\mathcal{B}}v_{1}(a,b)$ using the high-probability anytime regret bounds of $\texttt{ALG}_{2}$ as well as the upper bound on the Lipschitz constant. The algorithm then chooses the arm $a_{t}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a)$.
Initialize $\hat{v}_{1}(a)=1$ for $a\in\mathcal{A}$.
  // Initialize empirical means for $\max_{b\in\mathcal{B}}v_{1}(a,b)$.
Initialize $v^{\text{UCB}}_{1}(a)=1$ for $a\in\mathcal{A}$.
  // Initialize UCB for $\max_{b\in\mathcal{B}}v_{1}(a,b)$
@@
       Set $S(a):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},r)\in H\text{ s.t. }a=a_{t^{\prime}}\right\}$
        // Observed rewards
@@
             $\hat{v}_{1}(a)\leftarrow(\sum_{r\in S(a)}r)/|S(a)|$
              // Empirical mean
             $\alpha(a)\leftarrow\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{|S(a)|}}+C\cdot L\cdot\frac{\sqrt{\log T}}{\sqrt{|S(a)|}}$
              // confidence bound width
@@ 3             $v^{\text{UCB}}_{1}(a)\leftarrow\min(1,\hat{v}_{1}(a)+\alpha(a))$
Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\left(v^{\text{UCB}}_{1}(a)\right)$.
  // arm with max upper confidence bound
@@
Consider a StrongDSG where $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ has Lipschitz constant $L^{*}$. Let $\texttt{ALG}_{2}$ be any algorithm satisfying high-probability anytime regret $h(t,T,\mathcal{B})=C^{\prime}\sqrt{|\mathcal{B}|t\log T}$ where $C^{\prime}$ is a constant, and let $\texttt{ALG}_{1}=\texttt{LipschitzUCB}(L,C^{\prime}\sqrt{|\mathcal{B}|})$ for any $L\geq L^{*}$. Then both players achieve the following regret bounds with respect to the original Stackelberg benchmarks $\beta_{1}^{\text{orig}}$ and $\beta_{2}^{\text{orig}}$: that is, $R_{1}(T;\mathcal{I})=O\left(L\sqrt{T|\mathcal{A}||\mathcal{B}|\log T}\right)$ and $R_{2}(T;\mathcal{I})=O\left(L^{2}\sqrt{T|\mathcal{A}|\cdot|\mathcal{B}|\log T}\right)$.
@@
The intuition is the continuity conditions imply that small errors by the follower translate into bounded suboptimality for the leader (and vice versa); moreover, the high-probability anytime regret requirements bound the follower’s errors. Together, these properties guarantee that the leader’s empirical mean $\hat{v}_{1}(a)$ for each arm $a\in\mathcal{A}$ is close to the mean reward $v_{1}(a,b^{*}(a))$ that they would receive if the follower best-responded: in more detail, the main lemma (Lemma D.2) is that the empirical mean $\hat{v}_{1}(a)$ is at least $v_{1}(a,b^{*}(a))-\Theta(L\sqrt{\log T}/\sqrt{n_{t}(a)})$, where $n_{t}(a)$ is the number of times that arm $a$ has been pulled prior to time step $t$. Using Lemma D.2 to bound the suboptimality of the leader’s choice of actions $a_{t}\in\mathcal{A}$ and using the anytime regret requirements to bound the follower’s suboptimality, we can bound both the leader’s regret and the follower’s regret. We defer the full proof to Section D.2. ∎
@@ 5.2 Weaker benchmark
For the self-tolerant benchmarks, we show it is possible to achieve $\tilde{O}(\sqrt{T|\mathcal{A}||\mathcal{B}|})$ regret for both players (Theorem 5.3), which outperforms the $T^{2/3}$ lower bound for the stronger benchmark shown in Theorem 4.6. To demonstrate this is feasible, we focus on WeakDSGs, and we construct a specific pair of algorithms that achieve a $O(\sqrt{T})$ regret upper bound. For the follower, we take the algorithm $\texttt{ALG}_{2}$ to be ActiveArmElimination (Algorithm 7), which cycles through phases of exploration, after which all sufficiently suboptimal arms are eliminated. For the leader, we construct a UCB-based algorithm PhasedUCB (Algorithm 5) which constructs confidence bounds for every pair of actions $(a,b)$.
@@
The algorithm $\texttt{ALG}_{1}=\texttt{PhasedUCB}(M_{1},\ldots,M_{P})$ takes as input the parameters $M_{1},\ldots,M_{P}\geq 0$. (The parameter $M_{i}$ is intended to capture the number of times that an arm is pulled in phase $i$ by the instantiation of ActiveArmElimination specified by $\texttt{ALG}_{2}$.) The algorithm $\texttt{ALG}_{1}$ computes UCB estimates $v^{\text{UCB}}_{1}(a,b)$ for $v_{1}(a,b)$, computes the set of active arms $\mathcal{B}^{\prime}(a)$ in the previous phase of $\texttt{ALG}_{2}$’s instantiation of ActiveArmElimination for each arm $a$ (ComputeActiveArms, Algorithm 6), and chooses the arm with maximum UCB: $a_{t}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a,b)$. ComputeActiveArms computes the active arms $\mathcal{B}^{\prime}(a)$ by iterating through $H$ and keeping track of whenever a new phase is entered using the parameters $M_{1},\ldots,M_{P}$.
Let $\hat{v}_{1}(a,b)=0$ for $a\in\mathcal{A}$ and $b\in\mathcal{B}$.
  // initialize empirical mean of $v_{1}(a,b)$
Let $v^{\text{UCB}}_{1}(a,b)=1$ for $a\in\mathcal{A}$ and $b\in\mathcal{B}$.
  // initialize UCB for $v_{1}(a,b)$
Let $\mathcal{B}^{\prime}(a)=\texttt{ComputeActiveArms}(M_{1},\ldots,M_{P},H)$.
  // active arms in previous phase for $\texttt{ALG}_{2}$
@@
             Set $S(a,b):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},b_{t^{\prime}},r)\in H\text{ s.t. }a=a_{t^{\prime}},b=b_{t^{\prime}}\right\}$
              // observed rewards
@@
                   $\hat{v}_{1}(a,b)\leftarrow(\sum_{r\in S(a,b)}r)/|S(a,b)|$
                    // compute empirical mean
                   $\alpha(a,b):=10\cdot\sqrt{\frac{\log T}{|S(a,b)|}}$
                    // confidence bound width
                   $v^{\text{UCB}}_{1}(a,b)\leftarrow\min\left(1,\hat{v}_{1}(a,b)+\alpha(a,b)\right)$
                    // compute UCB
@@
Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\max_{b\in\mathcal{B}^{\prime}(a)}\left(v^{\text{UCB}}_{1}(a,b)\right)$.
  // arm with max upper confidence bound for any valid $b$
@@
Initialize $s^{\prime}(a)=0$ for $a\in\mathcal{A}$.
  // Index of the last completed phase for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $t^{\prime}(a)=1$ for $a\in\mathcal{A}$.
  // Time step marking beginning of phase $s^{\prime}+1$ for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $\mathcal{B}^{\prime}(a)=\mathcal{B}$.
  // Active arms in phase $s^{\prime}$ for $\texttt{ALG}_{2}$ on arm $a$.
Initialize $newphase_{a}=\text{False}$ for $a\in\mathcal{A}$.
  // Boolean for first time step in phase for $\texttt{ALG}_{2}$ on $a$.
@@
Consider a WeakDSG, where for each $a\in\mathcal{A}$, the algorithm $\texttt{ALG}_{2}$ runs a separate instantiation of ActiveArmElimination with parameters $M_{1},\ldots,M_{P}$ (where $M_{i}=\Theta(\log T\cdot 2^{2i})$ denotes the number of times that each arm is pulled in phase $i$). Let $\texttt{ALG}_{1}=\texttt{PhasedUCB}(M_{1},\ldots,M_{P})$. Then it holds that the regret with respect to the self-$\gamma$-tolerant benchmarks $\beta_{1}^{\text{self-tol}}$ and $\beta_{2}^{\text{self-tol}}$ is bounded as:
@@
The intuition is that the benchmark allows the leader to choose any $a\in\mathcal{A}_{\epsilon}$. The definition of $\mathcal{A}_{\epsilon}$ means that the leader can take an optimistic perspective on the follower’s choice of action $\mathcal{B}_{\epsilon}(a)$ (and not have to prepare for the worst-case action in $\mathcal{B}_{\epsilon}(a)$). This optimistic perspective surfaces in PhasedUCB in terms of how the leader evaluates an action $a$ based on the maximum UCB $\max_{b\in\mathcal{B}^{\prime}(a)}v^{\text{UCB}}_{1}(a,b)$ across all active arms $b\in\mathcal{B}^{\prime}(a)$. To analyze this pair of algorithms, we show a bound $\epsilon_{t}$ for each time step $t$ such that $a_{t}$ is an $\epsilon_{t}$-best-response for the leader and $b_{t}$ is an $\epsilon_{t}$-best-response for the follower: the main lemma (Lemma D.6) shows that we can set $\epsilon_{t}$ to be $\Theta(\sqrt{|\mathcal{B}\cdot\log T/n_{t}(a_{t})})$ where $n_{t}(a_{t})$ is the number of times that arm $a_{t}$ has been pulled prior to time step $t$. The full proof is deferred to Section D.3. ∎
@@
Consider StrongDSGs or WeakDSGs with actions sets $\mathcal{A}$ and $\mathcal{B}$ such that $|\mathcal{A}|\geq 2$ and $|\mathcal{B}|\geq 2$. For any algorithms $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$, there exists an instance $\mathcal{I}^{*}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ such that at least one of the players incurs $\Omega(\sqrt{T\cdot(\left|\mathcal{A}\right|-1)\cdot\left|\mathcal{B}\right|})$ regret with respect to the self-$\gamma$-tolerant benchmarks $\beta_{1}^{\text{self-tol}}$ and $\beta_{2}^{\text{self-tol}}$, that is: $\max(R_{1}(T;\mathcal{I}^{*}),R_{2}(T;\mathcal{I}^{*}))=\Omega(\sqrt{T\cdot(\left|\mathcal{A}\right|-1)\cdot\left|\mathcal{B}\right|})$.
@@
6 Discussion of Benchmark Parameters